OpenStudy (anonymous):

Solve this system of linear equations. -2x=8-6y -15x=21-6y

OpenStudy (anonymous):

you love me :D

OpenStudy (radar):

6y=2x+8 6y=15x+21 2x+8=15x+21 -13x=13 x=-1 Substitute that for x in first equation and solve for y.